Circular No. 8421

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 T3 An object discovered at Siding Spring on 0.5-m Uppsala Schmidt tele-

scope CCD images by R. H. McNaught (discovery observation below) has been found (after posting on the ‘NEO Confirmation Page’) to be diffuse on CCD images taken at Las Campanas with the 0.36-m ‘SoTIE’ reflector by G. Masi, F. Mallia, and R. Wilcox on Oct. 13, 17, and 21. Ten combined 1-min exposures taken on Oct. 13.1 UT show a slightly diffuse object with diameter ∼ 15′′ (slightly elongated northwest-southeast) with an apparent condensation in the northwest part of the coma; images taken on Oct. 17.3 show a possible 5′′ tail in p.a. 10^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Oct. 12.69016  3^{h}16^{m}05.^{s}77  −59^{o}40′24′′.5  18.7

The available astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2004-U24.

T = 2004 Nov. 11.633 TT  ω = 289.^{o}478
Ω = 47.909 }2000.0
q = 9.55852 AU  i = 72.760

COMET C/2004 U1 (LINEAR) An apparently asteroidal object discovered by LINEAR (discovery ob-

servation below) has been found to show a faint coma on CCD images taken on Oct. 20.1 UT by A. Knöfel (Schoenbrunn, Germany; 0.5-m f/5 reflector). CCD images taken by P. Birtwhistle (Great Shefford, U.K.; 0.30-m f/6.3 reflector) on Oct. 21.2 show the object to be possibly diffuse with diameter 9′′ and a possible broad tail 12′′ long in p.a. 320^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Oct. 19.41854  8^{h}56^{m}12.^{s}68  +20^{o}30′21′′.8  19.4

The available astrometry (including LONEOS prediscovery observations from Oct. 16), the following preliminary orbital elements, and an ephemeris appear on MPEC 2004-U25.

T = 2004 Dec. 4.319 TT  ω = 18.^{o}566
Ω = 112.608 }2000.0
q = 2.68042 AU  i = 130.288
2004 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green